It costs Crane Company $28 of variable costs and $10.00 of allocated fixed costs to produce an industrial trash can that sells for $50. A buyer in Mexico offers to purchase 3000 units at $30 each. Crane Company has excess capacity and can handle the additional production. What effect will acceptance of the offer have on net income? Increase $6000 Decrease $24000 Increase $24000 Increase $90000
Answer #1

Answer : Increase $ 6000

Variable cost per unit = $ 28

Offect Price Per unit = $ 30

Contribution MArgin per Unit on Offer = $30- 28 = $2

Order Size = 3000 Units

Total Contibution MArgin = 3000*2 = 6000

Increemntal Fixed cost = 0

Increase in Operating income = 6000-0 = 6,000